# NS-Net: Decoupling CLIP Semantic Information through NULL-Space for Generalizable AI-Generated Image Detection

### v1

@@ Introduction
Performance on GenImage. As shown in Table 1, our method demonstrates superior cross-domain generalization on GAN- and Diffusion-generated images, achieving an average accuracy of 93.2%. While its accuracy on real images (97.7%) is slightly lower than some methods (e.g., LadeDA at 99.8%), it surpasses the state-of-the-art AIDE (Yan et al. 2024) by 7.6% in mean accuracy (mAcc.). All methods achieve near-perfect detection (e.g., 100% on ProGAN and SDv1.4) on in-domain datasets; however, their performance degrades significantly on out-of-domain datasets. In contrast, our method maintains high accuracy across most cross-domain datasets, such as 100.0% on BigGAN and 98.9% on VQDM. This generalizability highlights the effectiveness of our approach in decoupling semantic information to focus on features critical for forgery detection.
@@
Performance on UniversalFakeDetect. Table 2 underscores our method’s capability in detecting unknown Diffusion-based fake images, achieving an average accuracy of 97.0%. It attains perfect detection (100%) on LDM_200, LDM_200_cfg, and DALLE, outperforming competing methods such as AIDE (95.6%) and DFFreq (92.9%). This consistent high performance across diverse diffusion models demonstrates the method’s strong generalization ability.
@@
Performance on AIGIBench. Given the rapid evolution of generative techniques, existing detection datasets often fall short in evaluating contemporary AI-generated image detection methods. We address this by utilizing the AIGIBench dataset (Li et al. 2025d), which includes a wide range of advanced generative models (Table 3). Our method still maintains the highest detection accuracy, with the improvement of 2.8% compared with the state-of-the-art methods. It significantly outperforms UnivFD by over 10% on most datasets, with a 21.4% improvement on GLIDE, owing to its ability to eliminate semantic information from CLIP’s upstream training while retaining fine-grained features essential for detection. However, performance on SocialRF (18.4%) and CommunityAI (8.9%) remains low, consistent with other methods, likely due to the heterogeneous and unknown generative methods in these datasets, suggesting further research.
Ablation Studies
Plug & Play Application In order to further verify the wide application capability of our method, we applied the NULL-Space Decoupling module and the Patch Slelection module to the existing detection framework. The results are shown in Table 7 and Table 8 of the Appendix. They all improve the detection ability of the model. More details can be found in Appendix IV.
Patch
Selection
Null-Space
Decoupling
Contrastive
Learning
mean Acc.
@@
Table 4: Ablation Study on the GenImage (Zhu et al. 2024) Dataset.
Conclusion
@@
Appendix A Appendix
Details of Detection Models for Comparison
@@ Appendix A Appendix
In order to effectively evaluate the advancement and generalization of our proposed method in the field of AI-generated detection, we selected several state-of-the-art detection methods from recent years. The following provide detailed descriptions of these detection approaches:
@@
1) CNN-Spot (CVPR 2020) (Wang et al. 2020). CNN-Spot uses CNN to identify synthetic content by analyzing common spatial artifacts in AI-generated images. It extracts hierarchical features from raw pixel data by stacking convolutional layers, effectively capturing generation anomalies.
@@
2) UnivFD (CVPR 2023) (Ojha, Li, and Lee 2023). UnivFD demonstrates that CLIP effectively extracts artifacts from images. By training a classifier on these features, they achieve strong generalization performance.
@@
3) FreqNet (AAAI 2024) (Tan et al. 2024a). FreqNet isolates high-frequency components of each image using an FFT-based high-pass filter, and introduces a plug-in frequency-domain learning block that transforms intermediate feature maps via FFT, applies learnable magnitude and phase transformations, and then performs an inverse FFT (iFFT), enabling optimization directly in the frequency domain.
@@
4) NPR (CVPR 2024) (Tan et al. 2024b). NPR targets the universal structural artifacts introduced by up-sampling layers in generative models. The method transforms each input image into NPR maps to capture signed intensity differences between each pixel and its four immediate neighbors. These maps make local pixel-dependency patterns explicit, revealing artifacts characteristic of synthetic up-sampling operations.
@@
5) Ladeda (arxiv 2024) (Cavia et al. 2024). LaDeDa is a patch-level deepfake detector that partitions each input image into 9 × 9 pixel patches and processes them using a BagNet-style ResNet-50 variant with its receptive field constrained to the same 9 × 9 region. The model assigns a deepfake likelihood to each patch, and the final prediction is obtained by globally pooling the patch-level scores.
@@
6) AIDE (ICLR 2025) (Yan et al. 2024). AIDE simultaneously incorporates low-level patch statistics and high-level semantics for AI-generated image detection. It employs two expert branches: i) a Semantic Feature Extractor, which utilizes CLIP-ConvNeXt embeddings to detect high-level content inconsistencies, and ii) a Patchwise Feature Extractor, which ranks image patches by spectral energy, selects the highest- and lowest-frequency regions, and applies a lightweight CNN to capture fine-grained noise and artifact patterns.
@@
7) DFFreq (arxiv 2025) (Yan et al. 2025). DFFreq first utilizea a sliding window to restrict the attention mechanism to a local window, and reconstruct the features within the window to model the relationships between neighboring internal elements within the local region. Then, it designs a dual frequency domain branch framework consisting of four frequency domain subbands of DWT and the phase part of FFT to enrich the extraction of local forgery features from different perspectives.
@@
8) SAFE (KDD 2025) (Li et al. 2025b). SAFE replaces conventional resizing with random cropping to better preserve high-frequency details, applies data augmentations such as Color-Jitter and RandomRotation to break correlations tied to color and layout, and introduces patch-level random masking to encourage the model to focus on localized regions where synthetic pixel correlations typically emerge.
@@
9) VIB-Net (CVPR 2025) (Zhang et al. 2025). VIB-Net finds that the general features extracted by current methods based on large-scale pre-trained models contain irrelevant features that are unrelated to the task of distinguishing real from fake images, and proposes VIB-Net, which uses Variational Information Bottlenecks to enforce authentication task-related feature learning.
Method Generative Adversarial Networks Other
ProGAN
CycleGAN
BigGAN
StyleGAN
StyleGAN2
GauGAN
StarGAN
WFIR
Deepfake
CNN-Spot 99.9 91.0 59.2 94.8 94.2 86.1 82.5 65.5 74.2
UnivFD 99.9 90.5 79.2 55.7 48.7 91.1 96.9 92.8 26.9
FreqNet 99.4 97.5 87.9 95.9 94.2 54.0 100.0 57.0 88.5
NPR 100.0 98.6 79.4 98.6 99.6 71.3 97.6 65.5 92.5
Ladeda 100.0 91.5 93.2 100.0 100.0 81.8 93.0 86.9 72.8
AIDE 99.6 98.2 91.6 99.4 99.9 74.1 99.7 90.8 66.5
DFFreq 99.7 99.0 98.4 100.0 100.0 98.1 99.8 89.0 87.5
SAFE 100.0 99.5 97.8 99.9 100.0 96.6 100.0 73.8 91.0
VIB-Net 99.9 99.7 97.1 93.9 98.1 99.2 99.7 95.7 83.6
Ours 100.0 100.0 99.7 99.8 99.9 97.0 100.0 99.5 95.2
Method Diffusion Models mA.P.
SDv1.4
SDv1.5
ADM
GLIDE
Midjourney
Wukong
VQDM
DALLE2
CNN-Spot 97.6 97.6 66.6 83.1 80.4 88.9 57.0 87.8 84.3
UnivFD 96.3 96.0 65.2 95.3 91.7 97.4 87.2 62.3 91.9
FreqNet 99.9 99.9 73.9 94.1 94.4 99.1 74.9 92.3 88.4
NPR 100.0 99.9 74.1 97.4 97.8 99.9 81.5 99.3 91.4
Ladeda 100.0 99.9 77.3 99.1 99.7 100.0 89.3 99.8 93.2
AIDE 100.0 99.9 99.4 99.9 90.3 99.9 99.3 99.9 94.5
DFFreq 100.0 99.9 98.7 99.5 99.7 100.0 99.3 99.9 98.1
SAFE 100.0 100.0 98.1 99.9 99.9 100.0 99.9 100.0 97.4
VIB-Net 100.0 99.9 91.2 96.4 94.3 99.9 97.4 94.6 96.5
Ours 100.0 100.0 100.0 99.8 97.0 100.0 100.0 100.0 99.2
Table 5: Cross-model Average Precision (A.P.) Performance on the GenImage (Zhu et al. 2024) Dataset.
Method Guided Glide_50_27 Glide_100_10 Glide_100_27 LDM_100 LDM_200 LDM_200_cfg DALLE mA.P.
CNN-Spot 70.8 76.9 78.0 76.8 88.7 90.1 93.8 62.7 79.7
UnivFD 81.0 92.5 92.1 92.3 91.4 91.4 84.7 80.3 88.2
FreqNet 78.5 97.8 97.9 97.6 99.7 99.8 99.7 93.1 95.5
NPR 83.0 99.3 99.0 99.3 99.9 99.9 100.0 98.9 97.4
Ladeda 89.0 99.6 99.4 99.7 99.9 99.9 99.9 96.4 98.0
AIDE 98.5 99.9 99.8 99.8 99.9 100.0 100.0 99.5 99.6
DFFreq 99.2 99.2 99.2 99.0 100.0 100.0 100.0 99.5 99.5
SAFE 98.9 99.6 99.7 99.4 100.0 100.0 100.0 99.9 99.7
VIB-Net 94.2 97.2 96.8 96.9 99.9 99.8 98.6 99.5 97.9
Ours 99.4 99.9 99.9 99.8 100.0 100.0 100.0 100.0 99.9
Table 6: Cross-model Average Precision (A.P.) Performance on the UniversalFakeDetect (Ojha, Li, and Lee 2023) Dataset.
Additional Metrics
In order to further study the effectiveness of our method and to be consistent with other research evaluation paradigms, we additionally use average precision as our evaluation metric. Average Precision (AP) provides a comprehensive evaluation of a model’s performance by integrating precision and recall across all decision thresholds, making it particularly effective for imbalanced datasets where the minority class is of primary interest. The specific results are shown in Table 5 and 6. Specifically, our method achieves the average precision of 99.2% and 99.9% on GenImage and UniversalFakeDetect respectively, surpassing the best methods 1.1% and 0.2%, further verifying the excellent detection effect and generalization ability of our proposed method.
Qualitative Analysis
@@
To further assess the generalization ability of our method, we visualize the feature distribution before binary classification, as shown in Table 3. This visualization highlights how effectively each trained model distinguishes between real and fake images, showcasing our method’s ability to generalize across diverse fake representations. From the visualization, it is evident that VIB-Net struggles with unseen GAN or diffusion models, as the feature distributions of real and generated images have a high overlap, indicating that the model often misclassifies fake images as "real". In contrast, our method demonstrates superior discrimination, effectively separating "real" and "fake" categories, even when encountering unseen sources.
@@
Table 7: Plug & Play Application of NULL-Space Decoupling on Existing Detectors. Cross-model Fake Accuracy (F.Acc.) Performance on the GenImage (Zhu et al. 2024) Dataset.
@@
Table 8: Plug & Play Application of Patch Selection on Existing Detectors. Cross-model Fake Accuracy (F.Acc.) Performance on the GenImage (Zhu et al. 2024) Dataset.
Plug & Play Application
@@
Patch Selection. We also performed a plug-and-play evaluation on our data preprocessing operation Parch Selection. The comparison results, illustrated in Table 8, demonstrate a consistent improvement in detecting synthetic images from various generative models, including the proposer of using crop during detection SAFE (Li et al. 2025b). The results demonstrate that Patch Selection enables the detector to extract more generalizable features indicative of artifacts, while mitigating the adverse effects of semantic content on artifact detection. This approach enhances the detector’s ability to identify subtle artifacts in input samples and improves its generalization performance.

### v2

@@ 4.2 Comparison with the State-of-the-Art
Performance on GenImage. As shown in Table 1 and Table 2, our method demonstrates superior cross-domain generalization on generated images, achieving 93.2% mAcc. and 99.2% mA.P. While its accuracy on real images (97.7%) is slightly lower than some methods, it surpasses the state-of-the-art AIDE [yan2024sanity] by 7.6% in mean accuracy (mAcc.). All methods achieve near-perfect detection (e.g., 100% on ProGAN and SDv1.4) on in-domain datasets; however, their performance degrades significantly on out-of-domain datasets. In contrast, our method maintains high accuracy across most cross-domain datasets, such as 100.0% on BigGAN and 98.9% on VQDM. This generalizability highlights the effectiveness of our approach in decoupling semantic information to focus on features critical for forgery detection.
@@
Performance on UniversalFakeDetect. Table 3 underscores our method’s capability in detecting unknown Diffusion-based fake images, achieving an average accuracy of 97.0%. It attains perfect detection (100%) on LDM and DALLE, outperforming competing methods such as AIDE (95.6%) and DFFreq (92.9%). This consistent high performance across diverse diffusion models demonstrates the method’s strong generalization ability.
@@
Performance on AIGIBench. Given the rapid evolution of generative techniques, existing detection datasets often fall short in evaluating contemporary AI-generated image detection methods. We address this by utilizing the AIGIBench dataset [li2025artificial], which includes a wide range of advanced generative models (Table 4). Our method still maintains the highest detection accuracy, with the improvement of 2.8% compared with the state-of-the-art methods. It significantly outperforms UnivFD by over 10% on most datasets, with a 21.4% improvement on GLIDE, owing to its ability to eliminate semantic information from CLIP’s upstream training while retaining fine-grained features essential for detection. However, performance on SocialRF (18.4%) and CommunityAI (8.9%) remains low, consistent with other methods, likely due to the heterogeneous and unknown generative methods in these datasets, suggesting further research.
Patch Selection Null-Space Decoupling Contrastive Learning mean Acc.
@@
Table 5: Ablation Study on the GenImage [zhu2024genimage] Dataset.
4.3 Ablation Studies
Plug & Play Application. To further verify the wide application capability of our method, we apply the NULL-Space Decoupling and Patch Selection to the existing detection framework. They all improve the detection ability of the model, as shown in Table 7 and Table 8 of the Appendix, suggesting that our method is more effective in promoting the extraction of more generalized artifact information.
4.4 Patch Selection Analysis
To further verify the effectiveness of our Patch Selection strategy, we conducted additional experiments, and the results are shown in Table 6. Using Resize or CenterCrop leads to a significant performance drop, while PatchShuffle can even perform worse than cropping on certain datasets, which applies additional shuffling after CenterCrop. Conversely, our Patch Selection yields consistent improvements across all datasets, indicating that it more comprehensively preserves artifact information. Furthermore, selecting only high-entropy or low-entropy patches fails to yield optimal performance, while RandomSelect achieves moderate gains, likely because it retains a mixture of both high- and low-entropy patches, further confirming their complementary roles in artifact preservation.
Selection Strategy GenImage [zhu2024genimage] AIGIBench [li2025artificial]
Resize 78.6 70.1
CenterCrop 89.1 75.9
PatchShuffle 87.3 76.9
Random 2k patches 88.4 81.2
Top 2k patches 85.8 75.1
Bottom 2k patches 90.1 78.8
Ours 93.2 87.0
Table 6: Selection Strategy Comparison on the GenImage [zhu2024genimage] and AIGIBench [li2025artificial] Datasets.
5 Conclusion
@@
[…]
Supplementary Material
6 Details of Detection Models for Comparison
@@
To effectively evaluate the advancement and generalization of our proposed method in AI-generated detection, we select several state-of-the-art detection methods from recent years:
@@
1) CNN-Spot (CVPR 2020) [wang2020cnn]. CNN-Spot uses CNN to identify synthetic content by analyzing common spatial artifacts in AI-generated images. It extracts hierarchical features from raw pixel data by stacking convolutional layers, effectively capturing generation anomalies.
@@
2) UnivFD (CVPR 2023) [ojha2023towards]. UnivFD demonstrates that CLIP effectively extracts artifacts from images. By training a classifier on these features, they achieve strong generalization performance.
@@
3) FreqNet (AAAI 2024) [tan2024frequency]. FreqNet isolates high-frequency components of each image using an FFT-based high-pass filter, and introduces a plug-in frequency-domain learning block that transforms intermediate feature maps via FFT, applies learnable magnitude and phase transformations, and then performs an inverse FFT (iFFT), enabling optimization directly in the frequency domain.
@@
4) NPR (CVPR 2024) [tan2024rethinking]. NPR targets the universal structural artifacts introduced by up-sampling layers in generative models. The method transforms each input image into NPR maps to capture signed intensity differences between each pixel and its four immediate neighbors. These maps make local pixel-dependency patterns explicit, revealing artifacts characteristic of synthetic up-sampling operations.
@@
5) Ladeda (arxiv 2024) [cavia2024real]. LaDeDa is a patch-level deepfake detector that partitions each input image into 9 × 9 pixel patches and processes them using a BagNet-style ResNet-50 variant with its receptive field constrained to the same 9 × 9 region. The model assigns a deepfake likelihood to each patch, and the final prediction is obtained by globally pooling the patch-level scores.
@@
6) AIDE (ICLR 2025) [yan2024sanity]. AIDE simultaneously incorporates low-level patch statistics and high-level semantics for AI-generated image detection. It employs two expert branches: a Semantic Feature Extractor, and a Patchwise Feature Extractor.
7) C2P-CLIP (AAAI 2025) [tan2025c2p]. C2P-CLIP concludes that CLIP achieves classification by matching similar concepts rather than discerning true and false. Based on this conclusion, they propose category common prompts to fine-tune the image encoder by manually constructing category concepts combined with contrastive learning.
@@
8) DFFreq (arxiv 2025) [yan2025generalizable]. DFFreq first utilizes a sliding window to restrict the attention mechanism to a local window, and reconstruct the features within the window to model the relationships between neighboring internal elements within the local region. Then, it designs a dual frequency domain branch framework consisting of four frequency domain subbands of DWT and the phase part of FFT to enrich the extraction of local forgery features from different perspectives.
@@
9) SAFE (KDD 2025) [li2025improving]. SAFE replaces conventional resizing with random cropping to better preserve high-frequency details, applies data augmentations such as Color-Jitter and RandomRotation to break correlations tied to color and layout, and introduces patch-level random masking to encourage the model to focus on localized regions where synthetic pixel correlations typically emerge.
@@
10) VIB-Net (CVPR 2025) [zhang2025towards]. VIB-Net finds that the general features extracted by current methods based on large-scale pre-trained models contain irrelevant features that are unrelated to the task of distinguishing real from fake images, and proposes VIB-Net, which uses VIB to enforce authentication task-related feature learning.
7 Qualitative Analysis
@@
To further assess the generalization ability of our method, we visualize the feature distribution before classification, as shown in Table 3. This visualization highlights how effectively each trained model distinguishes between real and fake images, showcasing our method’s ability to generalize across diverse fake representations. From the visualization, it is evident that VIB-Net struggles with unseen GAN or diffusion models, as the feature distributions of real and generated images have a high overlap, indicating that the model often misclassifies fake images as real. In contrast, our method demonstrates superior discrimination, effectively separating different categories, even when encountering unseen sources.
@@
Table 7: Plug & Play Application of NULL-Space Decoupling on Existing Detectors. Cross-model Fake Accuracy (F.Acc.) Performance on the GenImage [zhu2024genimage] Dataset.
@@
Table 8: Plug & Play Application of Patch Selection on Existing Detectors. Cross-model Fake Accuracy (F.Acc.) Performance on the GenImage [zhu2024genimage] Dataset.
8 Plug & Play Application
@@
Patch Selection. We also performed a plug-and-play evaluation on our data preprocessing operation Parch Selection. The comparison results, illustrated in Table 8, demonstrate a consistent improvement in detecting synthetic images from various generative models, including the proposer of using crop during detection SAFE [li2025improving]. The results demonstrate that Patch Selection enables the detector to extract more generalizable features indicative of artifacts, while mitigating the adverse effects of semantic content on artifact detection. This approach enhances the detector’s ability to identify subtle artifacts in input samples and improves its generalization performance.